|2 3 7 |
If |1 x -2 | = 5, find x.
|0 2 x |
1
Expert's answer
2019-07-09T12:37:30-0400

determinant is:

"2\u2217x\u2217x+3\u22170\u2217(\u22122)+7\u22171\u22172\u22127\u2217x\u22170\u22121\u22173\u2217x\u22122\u22172\u2217(\u22122)=5"

"2x^2 - 3x +17 = 0"

"\Delta = (-3)^2 - 4*2*17 = -127"

"\sqrt\Delta = \pm i\sqrt{127}"

"x = {3 \pm i\sqrt{127} \over 4}" or "x = {3 \over 4} \pm {\sqrt{127} \over 4}i"